# On $t$-structures adjacent and orthogonal to weight structures

### v1

@@ 2.3 Some properties and restrictions of strictly orthogonal $t$-structures
1. Let us formulate a simple general statement that we essentially used in the proof of Proposition 2.3.1(1); we will apply it below.
@@
2. In Definition 3.2 of [SaS11] a couple $(C_{1}^{\prime},C_{2}^{\prime})$ such that $C^{\prime}_{1},C^{\prime}_{2}\subset\operatorname{Obj}\underline{C}^{\prime}$, $C_{2}^{\prime}=C^{\prime}_{1}{}^{\perp}$, and $C^{\prime}_{1}={}^{\perp}C_{2}^{\prime}$, was said to be a Hom-orthogonal pair. It is well known that Hom-orthogonal pairs are closely related both to $t$-structures and to weight structures; see Definition 3.6 of ibid. and §3 of [BoV19].
Note here that the main object of study of ibid. are Hom-orthogonal pairs satisfying a certain "decomposition" axiom very similar to axioms (iv) in Definitions 1.2.1 and 1.1.3. They are said to be torsion theories; see Definition 2.1 of ibid.
@@
Assume that $\underline{C}^{\prime}\subset\underline{C}$, $\underline{D}=\underline{C}$, and $w$ is a weight structure on $\underline{C}$.
@@
(i). There exists a $t$-structure $t$ on $\underline{C}^{\prime}$ orthogonal to $w$.
@@
(iii). The functor $\tau_{\geq 0}H^{\underline{C}}_{M}=\tau_{\geq 0}H_{M}$ is $\underline{C}$-representable by an object of $\underline{C}^{\prime}$ for any object $M$ of $\underline{C}^{\prime}$ (that is, $\tau_{\geq 0}H_{M}\cong H_{L}$ for some object $L$ of $\underline{C}^{\prime}$).
(iv). If $M$ is an object of $\underline{C}^{\prime}$ then the functor $\tau_{\leq-1}H_{M}$ is isomorphic to $H_{R}$ for some $R\in\operatorname{Obj}\underline{C}^{\prime}$.
@@
Condition (ii) clearly implies condition (i). Conditions (ii) and (ii’) are equivalent according to Proposition 2.3.1. Next, Proposition 1.2.4(2) implies that $C_{2}\subset\underline{C}_{w\leq-1}$ (see (2.3.1) for the notation); hence $C_{1}\perp C_{2}$ by the definition of $C_{2}$. Hence Proposition 2.3.1 implies that $(i)\implies(ii)$.
Let us verify that condition (iv) implies (ii). So, we should check that the couple $(C_{2}[1],C_{1})$ is a $t$-structure. Now, these classes are clearly closed with respect to $\underline{C}^{\prime}$-isomorphisms, and the "shift" axiom (ii) of Definition 1.1.3 is obviously fulfilled as well. Moreover, we have just proved that $C_{1}\perp C_{2}$; this gives the orthogonality axiom (iii). Hence it remains to check the existence of $t$-decompositions (this is axiom (iv) of $t$-structures). For an object $M$ of $\underline{C}^{\prime}$ we assume $\tau_{\leq-1}H_{M}\cong H_{R}$ for some $R\in\operatorname{Obj}\underline{C}^{\prime}$. Take $\Phi$ to be the duality provided by Lemma 1.1.2. By the Yoneda lemma, the transformation $b_{-1}$ in (2.1.1) equals $\Phi(-,g)$ for some $g\in\underline{C}^{\prime}(M,R)$. Then Proposition 2.2.4(III) gives a $\underline{C}^{\prime}$-distinguished triangle $L\to M\to R\to L[1]$, where $L\in\operatorname{Obj}\underline{C}^{\prime}$ and $H_{L}\cong\tau_{\geq 0}H$. It remains to note that $L\in C_{1}$ and $R\in C_{2}$; see Proposition 2.2.2(1).
Next, (i) implies condition (v) according to Proposition 2.1.2(4). Moreover, conditions (v) clearly implies conditions (iii) and (iv). Lastly, assume that our condition (iii) is fulfilled; we fix $M\in\operatorname{Obj}\underline{C}^{\prime}$ and assume that $\tau_{\geq 0}H\cong H_{L}=\Phi({-,L})$ for some $L\in\operatorname{Obj}\underline{C}^{\prime}$. We argue similarly to the proof of (iv)$\implies$ (ii) above. By the Yoneda lemma, the corresponding transformation $a_{0}:\tau_{\geq 0}H\to L$ can be presented as $\Phi(-,f)$ for some $f\in\underline{C}^{\prime}(L,M)$. Then Proposition 2.2.4(III) yields that $\tau_{\leq-1}H_{M}\cong H_{R}$ for the corresponding $R$. Thus condition (iii) implies (iv), and this concludes the proof. II. If $M\in\underline{C}^{\prime}_{t=0}$ then the definition of orthogonality implies that the functor $H_{M}:\underline{C}^{op}\to\operatorname{Ab}$ (see Definition 2.1.1(3)) is of weight range $[0,0]$. Thus it suffices to apply Proposition 2.2.2(5) along with the Yoneda lemma and Lemma 2.1.4(1). ∎
Now we will describe certain restrictions of $t$ that is strictly ortogonal to $w$. We start from some notation.
@@
Throughout this paper $R$ will be an associative commutative unital ring. Assume that ${\underline{C}^{0}},\underline{C}\subset\underline{D}$ and ${\underline{C}^{0}}$ is $R$-linear.
@@
2. $\underline{C}_{\underline{A}\cap\underline{A}^{\prime}}=\underline{C}_{\underline{A}\cap\underline{A}^{\prime}}\cap\underline{C}_{\underline{A}\cap\underline{A}^{\prime}}$ (cf. Remark 1.1).
@@
2.4 On coproductive extensions of weight structures
@@
1. If $w$ is the coproductive extension of $w^{0}$ to $\underline{C}$ and $w^{\prime}$ is an arbitrary extension of $w^{0}$ to $\underline{C}$ then Proposition 1.2.4(2) is easily seen to imply that $\underline{C}_{w\leq 0}\subset\underline{C}_{w^{\prime}\leq 0}$; cf. Proposition 1.2.4(3). Moreover, if $w$ is an extension of $w^{0}$ to $\underline{C}$ such that $\underline{C}_{w\leq 0}\subset\underline{C}_{w^{\prime}\leq 0}$ for any other extension $w^{\prime}$ then this assumption determines it canonically (since $\underline{C}_{w\geq 0}\supset\underline{C}_{w^{\prime}\geq 0}$). Yet the existence of $w$ of this sort does not imply that $w$ is the coproductive extension of $w^{0}$ to $\underline{C}$ since the coproductive extension of $w^{0}$ to $\underline{C}$ does not havto o exist/ Indeed, the coproductive extension of $w^{0}$ does not exist in the case ${\underline{C}^{0}}=\{0\}$ and $\underline{C}\neq\{0\}$; yet $w=(0,\underline{C})$ clearly satisfies the aforementioned "minimality" property.
@@
Corollary 2.4.4 is a certain substitute for Theorem 2.2.5 of [Bon19]; see Remark 0.3 above. Though the author suspects that the results of the current text are not sufficient to prove loc. cit. itself, they can be successfully applied to treat all the examples of loc. cit. known to the author; cf. Proposition 3.3.5 below.
@@ 3 On $t$-structures orthogonal to (co)smashing weight structures
In §3.1 we consider smashing weight structures (on smashing triangulated categories); these are the ones that "respect coproducts".
@@
In §3.4 we formulate the dual to Theorem 0.1 and give some applications for it. If $w$ is both cosmashing and smashing then the left adjacent $t$-structure $t$ restricts to the subcategory of compact objects of $\underline{C}$ as well as to all other "levels of smallness" for objects. Combining this statement with an existence of weight structures theorem from [KeN13] we obtain a statement on $t$-structures extending yet another result of ibid. We will not apply these results in the current paper.
@@ 3.1 On smashing weight structures and related notions
We will need a few definitions.
@@ 3.2 On the existence of $t$-structures adjacent to smashing weight structures
    We will say that $\underline{C}$ satisfies the Brown representability property whenever any cp functor from $\underline{C}$ into $\operatorname{Ab}$ is representable.
@@
    We will say that a generating class of objects $\mathcal{P}$ as well as the subcategory $C$ of $\underline{C}$ with $\operatorname{Obj}C=\mathcal{P}$ compactly generate $\underline{C}$ if $\mathcal{P}$ is also essentially small and consists of compact objects (of $\underline{C}$).
@@
Now we prove our first "practical" existence of $t$-structures results; see Definitions 3.1.1 and 3.2.1 for the notions mentioned in our theorem.
@@
Then there exists a $t$-structure $t^{r}$ left adjacent to $w$ if and only if $w$ is smashing. Moreover, $t^{r}$ is cosmashing (if exists) and its heart is equivalent to the category of those additive functors ${\underline{Hw}}^{op}\to\operatorname{Ab}$ that respect products.
@@
The "only if" assertion is essentially given by Proposition 2.4(6) of [BoV19] (cf. §3. of ibid.; the statement is also very easy for itself).
Conversely, assume that $w$ is smashing. According to Theorem 2.3.3(I) the existence of $t^{r}$ is equivalent to the representability of $\tau_{\leq 0}H_{M}$ for any representable functor $H_{M}$. Next, Proposition 3.1.2(5) says that $\tau_{\leq 0}H_{M}$ is a cp functor since $H_{M}$ is. Hence $\tau_{\leq 0}H_{M}$ is representable by the Brown representability assumption, and we obtain that $t^{r}$ exists indeed.
Next, the category $\underline{C}$ is cosmashing according to Proposition 8.4.6 of [Nee01] (since it satisfies the Brown representability property). Moreover, $t^{r}=(C_{2}[1],C_{1})$, where $C_{1}=\underline{C}_{w\leq-1}^{\perp}$ and $C_{2}=\underline{C}_{w\geq 0}^{\perp}$ according to Theorem 2.3.3(I) (alternatively, one can apply Proposition 1.3.3). Hence the class $\underline{C}_{t^{r}\geq 0}$ is closed with respect to $\underline{C}$-products; thus $t^{r}$ is cosmashing as well.
Lastly, since $t^{r}=(C_{2}[1],C_{1})$, the class $\underline{C}_{t=0}$ equals $(\underline{C}_{w\geq 1}\cup\underline{C}_{w\leq-1})^{\perp}$; hence ${\underline{Ht}}$ can be calculated using Proposition 2.3.2(8) of [Bon21] (see also Remark 2.1.3(2) of ibid. and Proposition 2.2.2(5) above; cf. Theorem 2.3.3(II)). ∎
@@
1. Now let us discuss examples to Theorem 3.2.3.
According to Theorem 5 of [Pau12], any set $\mathcal{P}$ of compact objects of $\underline{C}$ generates a (unique) smashing weight structure (see Definition 3.2.1(5) and Remark 3.2.2(2)). Moreover, Theorem 4.5(2) of [PoS16] and Theorem 3.2.1 of [Bon22a] (we will mention these statements in the proof of Corollary 3.4.3(2) below) enable one to check whether two weight structures obtained this way are distinct. Thus one may say that there are lots of smashing weight structures on $\underline{C}$ whenever there are "plenty" of compact objects in it (see Theorem 4.15 of [PoS16] for a certain justification of this claim for derived categories of commutative rings). Thus our theorem yields a rich collection of $t$-structures, and the author does not know of any other methods that give all of them (cf. Remark 0.2).
Other notable statements related to the construction of smashing weight structures are Theorem 2.3.4(3) of [Bon22a] (it says that any perfect set of objects in a smashing triangulated category generates a smashing weight structure) and Theorem 3.1.3 and 3.2.2 of [BoS19] that treat weight-exact localization functors.
2. Now let us apply our theorem in the case of semi-orthogonal decompositions (see Definition 1.2.5 and Remark 1.2.6).
If $w$ is a semi-orthogonal decomposition of $\underline{C}$ (where $\underline{C}$ satisfies the Brown representability property) that is smashing as a weight structure then for the corresponding $t^{r}$ the couple $(\underline{C}_{t^{r}\geq 0},\underline{C}_{t^{r}\leq 0})$ is a semi-orthogonal decomposition as well; see Proposition 1.3.5. Note also that $(\underline{C}_{t^{r}\geq 0},\underline{C}_{t^{r}\leq 0})$ is a (cosmashing) weight structure.
Thus if we apply Proposition 1.3.5 to the semi-orthogonal decomposition $(\underline{C}_{t^{r}\geq 0},\underline{C}_{t^{r}\leq 0})$ we obtain that the embedding $\underline{C}_{2}\to\underline{C}$ possesses a right adjoint as well; note that $\underline{C}_{t^{r}\geq 0}=\underline{C}_{w\geq 0}$. Thus $\underline{C}_{2}$ is admissible in $\underline{C}$ in the sense of [BoK89, Definition 2.5] and the embedding $\underline{C}^{\prime}\to\underline{C}$ may be completed to a gluing datum (cf. [BBD82, §1.4] or [Nee01, §9.2]).

### v2

@@ 2.3 Strictly orthogonal $t$-structures (and their restrictions)
Let us formulate a simple general statement that we essentially used in the proof of Proposition 2.3.1(1); we will apply it below.
@@
Assume that $\underline{C}^{\prime}\subset\underline{C}$, and $w$ is a weight structure on $\underline{C}$.
@@
(i). There exists a $t$-structure $t$ on $\underline{C}^{\prime}$ orthogonal to $w$ (in $\underline{C}$).
@@
(iii). The functor $\tau_{\geq 0}H^{\underline{C}}_{M}=\tau_{\geq 0}H_{M}$ is $\underline{C}$-representable by an object of $\underline{C}^{\prime}$ for any object $M$ of $\underline{C}^{\prime}$ (that is, $\tau_{\geq 0}H_{M}\cong H_{LM}$ for some object $LM$ of $\underline{C}^{\prime}$).
(iv). If $M$ is an object of $\underline{C}^{\prime}$ then the functor $\tau_{\leq-1}H_{M}$ is isomorphic to $H_{RM}$ for some $RM\in\operatorname{Obj}\underline{C}^{\prime}$.
@@
Condition (ii) clearly implies condition (i). Conditions (ii) and (ii’) are equivalent according to Proposition 2.3.1. Next, Proposition 1.2.4(2) implies that the class $C_{2}$ (see (2.3.1) for this notation) lies in $\underline{C}_{w\leq-1}$; hence $C_{1}\perp C_{2}$ by the definition of $C_{2}$. Hence Proposition 2.3.1 implies that $(i)\implies(ii)$.
Let us verify that condition (iv) implies (ii). So, we should check that the couple $(C_{2}[1],C_{1})$ is a $t$-structure. Now, these classes are clearly closed with respect to $\underline{C}^{\prime}$-isomorphisms, and the "shift" axiom (ii) of Definition 1.1.1 is obviously fulfilled as well. Moreover, we have just proved that $C_{1}\perp C_{2}$; this gives the orthogonality axiom (iii). Hence it remains to check the existence of $t$-decompositions (this is axiom (iv) of $t$-structures). For an object $M$ of $\underline{C}^{\prime}$ we assume $\tau_{\leq-1}H_{M}\cong H_{RM}$ for some $RM\in\operatorname{Obj}\underline{C}^{\prime}$. By the Yoneda lemma, the transformation $b_{-1}$ in (2.1.1) equals $\underline{C}(-,g)$ for some $g\in\underline{C}^{\prime}(M,RM)$. Then Proposition 2.2.4(III) gives a $\underline{C}^{\prime}$-distinguished triangle
$LM\to M\to RM\to LM[1],$
where $LM\in\operatorname{Obj}\underline{C}^{\prime}$ and $H_{LM}\cong\tau_{\geq 0}H$. It remains to note that $LM\in C_{1}$ and $RM\in C_{2}$; see Proposition 2.2.2(1).
Next, (i) implies condition (v) according to Proposition 2.1.2(4). Moreover, condition (v) clearly implies conditions (iii) and (iv).
Lastly, assume that our condition (iii) is fulfilled; we fix $M\in\operatorname{Obj}\underline{C}^{\prime}$ and assume that $\tau_{\geq 0}H\cong H_{LM}=\underline{C}({-,LM})$ for some $LM\in\operatorname{Obj}\underline{C}^{\prime}$. We argue similarly to the proof of ((iv)$\implies$ (ii)) above. By the Yoneda lemma, the corresponding transformation $a_{0}:\tau_{\geq 0}H\to H_{LM}$ can be presented as $\underline{C}(-,f)$ for some $f\in\underline{C}^{\prime}(LM,M)$. Then Proposition 2.2.4(III) yields that $\tau_{\leq-1}H_{M}\cong H_{RM}$ for the corresponding $RM$. Thus condition (iii) implies (iv), and this concludes the proof.
II. If $M\in\underline{C}^{\prime}_{t=0}$ then the definition of orthogonality implies that the functor $H_{M}:\underline{C}^{op}\to\operatorname{Ab}$ (see Definition 2.1.1(3)) is of weight range $[0,0]$. Thus it suffices to apply Proposition 2.2.2(5) along with the Yoneda lemma and Lemma 2.1.4(1). ∎
Now we will describe certain restrictions of $t$ that is strictly orthogonal to $w$. We start from some definitions.
@@
Assume that ${\underline{C}^{0}},\underline{C}\subset\underline{D}$, and ${\underline{C}^{0}}$ is $R$-linear.
@@
2. We will say that $\underline{A}^{0}\subset\underline{A}$ is an abelian subcategory of $\underline{A}$ if $\underline{A}$ is its (strictly full) subcategory that contains the $\underline{A}$-kernel and the $\underline{A}$-cokernel of any $\underline{A}^{0}$-morphism.
@@
2. $\underline{C}_{\underline{A}\cap\underline{A}^{\prime}}=\underline{C}_{\underline{A}\cap\underline{A}^{\prime}}\cap\underline{C}_{\underline{A}\cap\underline{A}^{\prime}}$.
@@
2.4 Coproductive extensions of weight structures
@@
1. If $w$ is the coproductive extension of $w^{0}$ to $\underline{C}$ and $w^{\prime}$ is an arbitrary extension of $w^{0}$ to $\underline{C}$ then Proposition 1.2.4(2) is easily seen to imply that $\underline{C}_{w\leq 0}\subset\underline{C}_{w^{\prime}\leq 0}$; cf. Proposition 1.2.4(3). Moreover, if $w$ is an extension of $w^{0}$ to $\underline{C}$ such that $\underline{C}_{w\leq 0}\subset\underline{C}_{w^{\prime}\leq 0}$ for any other extension $w^{\prime}$ then this assumption determines it canonically (since $\underline{C}_{w\geq 0}\supset\underline{C}_{w^{\prime}\geq 0}$). Yet the existence of $w$ of this sort does not imply that $w$ is the coproductive extension of $w^{0}$ to $\underline{C}$ since the coproductive extension of $w^{0}$ to $\underline{C}$ does not have to exist. Indeed, the coproductive extension of $w^{0}$ does not exist in the case ${\underline{C}^{0}}=\{0\}$ and $\underline{C}\neq\{0\}$; yet $w=(0,\underline{C})$ clearly satisfies the aforementioned minimality property.
@@
Corollary 2.4.4 is a certain substitute for Theorem 2.2.5 of [Bon19]; see Remark 6.3 below. Though the author suspects that the results of the current text are not sufficient to prove loc. cit. itself, they can be successfully applied to treat all the examples of loc. cit. known to the author; cf. Proposition 3.3.5 below.
@@ 3 On $t$-structures orthogonal to (co)smashing weight structures
In §3.1 we consider smashing weight structures (on smashing triangulated categories); these are the ones that ’respect coproducts’.
@@
In §3.4 we formulate the dual to Theorem 0.1 and give some applications for it. If $w$ is both cosmashing and smashing then the left adjacent $t$-structure $t$ restricts to the subcategory of compact objects of $\underline{C}$ as well as to all other ’levels of smallness’ for objects. Combining this statement with an existence of weight structures theorem from [KeN13] we obtain a statement on $t$-structures extending yet another result of ibid.
@@ 3.1 On smashing weight structures and related notions
We need some definitions related to (co)products.
@@ 3.2 On the existence of $t$-structures adjacent to smashing weight structures
    We will say that $\underline{C}$ satisfies the Brown representability property whenever any cp functor (see Definition 2.4.1(2)) from $\underline{C}$ into $\operatorname{Ab}$ is representable.
@@
    We will say that a generating class of objects $\mathcal{P}$ (as well as the subcategory $C$ of $\underline{C}$ with $\operatorname{Obj}C=\mathcal{P}$) compactly generates $\underline{C}$ if $\mathcal{P}$ is also essentially small and consists of compact objects (of $\underline{C}$).
@@
Now we prove our first ’practical’ existence of $t$-structures result; see Definitions 3.1.1 and 3.2.1 for the notions mentioned in our theorem.
@@
Then there exists a $t$-structure $t^{l}$ left adjacent to $w$ if and only if $w$ is smashing. Moreover, $t^{l}$ is cosmashing (if exists) and its heart is equivalent to the category of those additive functors ${\underline{Hw}}^{op}\to\operatorname{Ab}$ that respect products.
@@
The "only if" assertion is essentially given by Proposition 2.4(6) of [BoV19] (cf. §3 of ibid.; the statement is also very easy for itself).
Conversely, assume that $w$ is smashing. According to Theorem 2.3.3(I) the existence of $t^{l}$ is equivalent to the representability of $\tau_{\leq 0}H_{M}$ for any representable functor $H_{M}$. Next, Proposition 3.1.2(5) says that $\tau_{\leq 0}H_{M}$ is a cp functor since $H_{M}$ is. Hence $\tau_{\leq 0}H_{M}$ is representable by the Brown representability assumption, and we obtain that $t^{l}$ exists indeed.
Next, the category $\underline{C}$ is cosmashing according to Proposition 8.4.6 of [Nee01] (since it satisfies the Brown representability property). Moreover, $t^{l}=(C_{2}[1],C_{1})$, where $C_{1}=\underline{C}_{w\leq-1}^{\perp}$ and $C_{2}=\underline{C}_{w\geq 0}^{\perp}$ according to Theorem 2.3.3(I) (alternatively, one can apply Proposition 1.3.3). Hence the class $\underline{C}_{t^{l}\geq 0}$ is closed with respect to $\underline{C}$-products; thus $t^{l}$ is cosmashing as well.
Lastly, since $t^{l}=(C_{2}[1],C_{1})$, the class $\underline{C}_{t^{l}=0}$ equals $(\underline{C}_{w\geq 1}\cup\underline{C}_{w\leq-1})^{\perp}$. Hence Proposition 2.2.2(5,10) along with Proposition 2.3.2(8) of [Bon21] gives the calculation in question. ∎
@@
Now let us discuss examples to Theorem 3.2.3.
1. According to Theorem 5 of [Pau12], any set $\mathcal{P}$ of compact objects of $\underline{C}$ generates a (unique) smashing weight structure (see Definition 3.2.1(5) and Remark 3.2.2(2)). Thus one may say that there are lots of smashing weight structures on $\underline{C}$ whenever there are ’plenty’ of compact objects in it (see Theorem 4.15 of [PoS16] for a certain justification of this claim for derived categories of commutative rings). Thus our theorem yields a rich collection of $t$-structures; yet cf. Remark 6.4 below.
We will say more on smashing weight structures in Remark 6.5 below. Moreover, we will also discuss the application of Theorem 3.2.3 to semi-orthogonal decompositions; this gives a generalization of [NiS09, Corollary 2.4].
2. The author has applied Theorem 3.2.3 in §4.2 of [BoK21] to obtain and study an interesting family of $t$-structures on various motivic triangulated categories.